# breaklines

### v3

@@ 5.1. Type Translation
Automatically resolving types is a challenging problem (Terekhov and Verhoef, 2000; Guizzo et al., 2023), and a large body of work has attempted to address this, mostly using symbolic rule-based approaches (Qiu, 1999; Dony et al., 2010; Broggi and Liu, 2023; Roziere et al., 2020; Eman J. Coco and Osman, 2018; Lano and Siala, 2024). AlphaTrans employs a Retrieval-Augment Generation (RAG) (Lewis et al., 2020) technique for finding the equivalent types in target language. To that end, it first extracts all the types in the source language of a given project. Custom, application-level types will be resolved during the translation as AlphaTrans translates fragments and classes in the target language. For the remaining types, it crawls the API documentation of the source language and retrieves the relevant description of each type.
To form the prompt [^3] Due to space limit, we could not include the prompt. Please refer to the artifacts to see the prompts used for type resolution., AlphaTrans uses the retrieved description and instructs the model with an in-context example to return the most relevant type in the target language, given the use of types in the source language and the retrieved description. To account for potential hallucination in LLM’s response, i.e., returning a type that does not exist in Python, AlphaTrans employs a simple Python script, uses the translated type as an annotation, executes the script, and keeps those without any syntactic or runtime issue. The types in the source language and their corresponding in the target language form a data structure called Universal Type Mapping. In practice, AlphaTrans reuses or augments the mapping when translating new projects.
@@ 5.2. Skeleton Construction
At the first step, AlphaTrans creates classes corresponding to each application class in Java. The fields for Python classes are set to None, and AlphaTrans uses the information in schema (§4.2.1) to ensure the naming corresponds to the type of their access modifier in the source language. In the example of Figure 4, the translation of field private Option option; in Java is __option: Option = None. The classes also include method signatures, with their body set to pass. AlphaTrans uses the universal type mapping (§5.1) to use relevant types in the full method signature. Once the initial skeleton is created, AlphaTrans leverages the extracted call graph during program decomposition to detect circular dependencies. If there exist any, AlphaTrans resolves them with local imports. For inner classes, AlphaTrans unfolds them in Python and uses dot notation to access specific methods and fields (e.g., Class.methodName). Finally, AlphaTrans implements best practices in Python and sub-classes all abstract classes and interfaces from abc.ABC class. Here, ABC is a class from the abc module in the Python standard library, which is used for defining abstract base classes.
@@ 6. Compositional Translation and Validation
AlphaTrans translates method fragments in the reverse call order. It takes the project’s call graph, removes the back edges to make it acyclic, computes topological order (i.e., linear ordering of its vertices), and translates method fragments (corresponding to vertices) in reverse topological order. Field fragments are independent; therefore, AlphaTrans translates them before method fragments. The algorithm takes a fragment $F$, LLM $M$, Skeleton $S,$ and a series of parameters as inputs, translates the fragment, recomposes the skeleton with the successfully translated fragment, and returns translation outcomes: (1) syntax check (denoted by the “non-parseable” and “parseable” labels), (2) functional equivalence check (denoted by the “graal-fail”, “graal-success”, and “graal-error” labels), and (3) translated test execution check (denoted by “not-exercised”, “test-fail”, and “test-success” labels).
AlphaTrans employs iterative and feedback-based prompting. That is, if one of the mentioned checks fails, e.g., the translated fragment is not syntactically correct, it prompts the model for another translation attempt. To control the number of iterations, AlphaTrans considers a reprompting budget, i.e., repromptBudget. The algorithm takes the minimum ($min_{budget}$) and maximum ($max_{budget}$) values for the budget and dynamically sets reprompting budget to a number between them based on the coverage information, e.g., the budget is close to $max_{budget}$ for a fragment if it is exercised multiple times (high hit rate based on coverage information) by different unit tests. The rationale is to give more importance to fragments covered by more tests to eventually increase translation validation success. The main translation and validation loop (lines 3–31) runs until the assigned budget is exhausted.
@@ 44       $break$;
Inside the loop, AlphaTrans first crafts a unique prompt based on the template shown in Figure 5 and then instructs the LLM to translate the fragment (lines 4–5). It then validates the generated translation in multiple steps. The first step checks for syntactical correctness and assigns proper labels to $TVO$ (lines 6–10). Then, AlphaTrans leverages GraalVM for isolation-based validation of fragment $F$ (lines 12–20), if there exists a test in the source language covering the fragment during its execution.
Finally, it translates and executes decomposed fragment tests: if there are no eligible tests (a test becomes eligible if all its dependencies are translated) for the fragment, AlphaTrans simply assigns the “not-exercised” label to the fragment and moves on to the next one (lines 21–23). Otherwise, it translates the tests, executes them to validate the fragment, and assigns test outcome labels in TVO (lines 24–31). In case of a test failure, AlphaTrans extracts all involved fragments and re-prompts them with feedback extracted from test execution.
Due to inherent intra- and inter-procedural dependencies in real-world projects, the number of fragments involved in re-prompting could be high, logarithmically increasing the translation time. AlphaTrans filters out those with GraalVM label “graal-success”, ranks the remaining based on suspiciousness score, and re-prompts $topK$ suspicious fragments. The suspiciousness score for fragments is calculated such that a fragment with more failing tests will get a higher score and, therefore, ranked higher among other fragments.
Our prompt template consists of five distinct parts (Figure 5). The first part is the persona message used by DeepSeek-Coder-33b-Instruct during instruction fine-tuning and is required for producing the best outputs. The next part introduces the In-Context Learning (ICL) example, which reflects the complexities of code translation and instructs LLM on how to deal with them. The green part indicates the natural language instruction given to the model. After describing the objective, the prompt embeds the source Java code along with partial translation as a skeleton, which includes all dependencies and translations of the fragments invoked by the current one. The prompt concludes with ### Response: special keyword to guide the model for code generation.
@@ 6.1. Language Interoperability
GraalVM (Würthinger et al., 2013; Oracle, 2024) is a Java Development Kit by Oracle. It offers the Polyglot API (GraalVM, 2024), which allows the integration of programs written in different guest languages within a Java-based host application. In the context of this paper, GraalVM allows developers to execute Python code from Java and vice versa. AlphaTrans leverages the Polyglot API to perform in-isolation validation of the fragments by replacing the Java implementation of a method with its translated Python version while keeping the rest of the project in Java. It then executes the Java tests covering the fragment to validate the functional equivalence of the translation.
The Polyglot API allows access to Python data objects from Java and vice-versa, as these objects reside in a shared memory space. However, objects must be cast to appropriate types for passing parameters to and processing returned values from polyglot calls. The Polyglot API can perform this casting implicitly for only a few simple data types. AlphaTrans builds on top of the Polyglot API to provide a framework to create a Python program state that is isomorphic to the Java program state. The Python translation is restricted to this isomorphic state, and the states are synchronized after method calls to preserve the isomorphism. AlphaTrans allows for the casting of user-defined types as well as several built-in and library types. Using both the static and the dynamic type information of Java objects, AlphaTrans can disambiguate the target types when casting Python objects to Java types. It further preserves object identities and aliasing during such casting and propagates exceptions across language boundaries.
@@ 6.3. Test Translation
Similar to translating application fragments, AlphaTrans also translates test fragments. Using the dependency information captured during static analysis, it crafts prompts for unit tests along with their dependencies for the model to translate. The ICL examples used for prompting test fragments differ from prompts used for translating application method fragments. The focus of ICL examples here is to prevent LLM from hallucinating the usage of assert statements in the source to target PL. To construct ICL examples for test fragment translation, we created a pool of in-context examples, where each example shows the Python assert statements equivalent to Java assert statements in the context of a test. When prompting a test fragment, the AlphaTrans detects the assert statement in the fragment and retrieves the corresponding examples from the pool. For translated tests, only syntactic validation is performed as there is no other means of validating their translations.
@@ 1. RQ1:
    Effectiveness. To what extent AlphaTrans can automatically resolve types from source to target PL? Can AlphaTrans effectively translate real-world projects?
@@ 7.1. Experiment Setup
AlphaTrans uses CodeQL (GitHub, 2024) and tree-sitter (Tree-Sitter, 2024) for static analysis. For running tests, validating translation, and computing coverage, AlphaTrans uses GraalVM $21.0.3+7.1$ (Oracle, 2024), JUnit 4 and 5 (Team, 2024b), Pytest 8.2.1 (pytest dev, 2024), JaCoCo (Team, 2024a), and Python’s coverage (Batchelder, 2024). AlphaTrans works with API- and open-access LLMs. We considered the following criteria for selecting the LLM: (1) for better reproducibility, we prioritized open-access models; (2) due to computing constraints, we wanted an LLM with moderate size ($<10B$ but $<70B$ parameters); (3) the model should perform reasonably well in code-related tasks; and (4) the model should have fast inference time due to the huge number of prompts. Per the mentioned criteria, we selected DeepSeek-Coder-33b-Instruct for main experiments (RQ1–RQ4). We also used GPT-4o, one of the best-performing commercial models, in RQ5 (§7.6) to demonstrate the impact of stronger models on improving the performance of AlphaTrans. We prompted models under the temperature 00 setting for reproducibility and used their default setting for other parameters. For the base prompting (Algorithm 2), we set the minimum and maximum values of the reprompting budget to $3$ and $5$. For the feedback prompting, we set the reprompting budget to $1$, i.e., AlphaTrans attempts to fix issues with feedback only once.
@@ 7.2.1. Effectiveness in type resolution and skeleton validation:
AlphaTrans extracted $1,797$ distinct types from the source projects and attempted to translate them to equivalent Python types. Of these, $915$ are application-level types (e.g., classes defined within the Java projects) and were directly resolved during skeleton construction. AlphaTrans prompts DeepSeek-Coder-33b-Instruct to resolve the remaining $882$, out of which it successfully translated $738~{}((915+738)/1,797=91.99\%)$ of them: generated types passed the syntactic and runtime check. The column ATR in Table 1 shows the results of automated type resolution. Since type resolution is essential to project skeleton construction, we manually checked the type mappings generated by DeepSeek-Coder-33b-Instruct, and also attempted to translate the remaining unresolved $144$ types.
Through manual investigation of the automatically resolved types, we observed that DeepSeek-Coder-33b-Instruct’s type resolution for $182$ cases, while correct, can be improved. For example, AlphaTrans translated java.io.File, a class concerning file manipulation functionality to str. The resolved type can represent file paths in Python but lacks features for file manipulation. We suspect this translation is impacted by the Java use case provided in the prompt. While this translation is correct concerning the use case, we replaced it with pathlib.Path to have a more generic type mapping. We also augmented the type mapping with additional types in the target language for $38$ types. For example, AlphaTrans translated java.nio.Buffer to bytearray, which is correct as they both provide a mutable sequence of bytes with efficient in-place modifications. However, array.array and memoryview also provide similar functionality with efficient and low-level data manipulation capabilities. Consequently, we augmented type mapping to typing.Union[bytearray, array.array, memoryview]. Given that type mapping can be reused, this one-time manual effort increases the chance of AlphaTrans’s success on unseen projects.
@@
Using the universal type mapping, AlphaTrans successfully creates and validates project skeletons in target PL, achieving $100\%$ syntax and runtime validation (column SV in Table 1). The skeleton validation step ensures all module imports, class structures, method signatures, and type annotations are done properly, making the subsequent steps easier. Applying AlphaTrans to unseen projects, if a class skeleton cannot be validated, AlphaTrans removes it from the target project, updates the skeleton based on the class dependencies, and proceeds to the next phase.
@@ 7.2.2. Effectiveness in compositional translation and validation:
Table 2 shows the compositional translation and validation results. The AMF column indicates the total number of application method fragments. The numbers in subsequent columns demonstrate the effectiveness of AlphaTrans in the translation and validation of AMFs only [^4] Due to space limit, please refer to our artifact (Lab, 2025) for details of all the translation and validation of other fragments shown in Table 1.. The Syntax Check column indicates the percentage of AMFs that pass syntactic validation. Column SNEF shows the percentage of AMFs not covered by JUnit tests. AlphaTrans successfully generates syntactically correct code ($98.80\%$ of AMFs and $96.40\%$ of all fragments—field, application, and test fragments—across ten subjects). We also observe that $43.43\%$ of AMFs are not covered during the execution of any test, i.e., we cannot go beyond syntactic check and validate their runtime behavior or functional equivalence.
For $56.57\%$ of AMFs that can be covered by source project tests, AlphaTrans attempts to validate the functional equivalence using GraalVM. The super column GraalVM shows the percentage of AMFs that GraalVM executes and successfully validates (GS), executes but there is a test assertion failure (GF), and cannot execute due to its limitation (GE) mentioned in §6.1. $24.50\%$ (min=$6.01\%$ and max=$70.70\%$), $16.23\%$ (min=$1.04\%$ and max=$32.50\%$), and $15.84\%$ (min=$3.13\%$ and max=$25.11\%$) of AMFs resulted in Graal Success, Graal Fail, and Graal Error, respectively. Note that these numbers add up to $56.57\%$ of AMFs that were covered by source program tests. With respect to only covered methods, GraalVM Success will be $43.30\%$. Furthermore, our analysis shows that a high portion of methods that are not covered by tests are either abstract methods or getter/setter methods. If their translations are syntactically correct, they are also likely functionally equivalent, which can ramp up the success rate (§7.5). Spearman Rank Order Correlation (Spearman, 1961) indicates a strong positive correlation between the method coverage (Table 1) and GS numbers ($\rho=0.92$), confirming that with better method coverage, the validated AMFs are very likely to be higher.
@@
Regardless of GraalVM’s validation for functional equivalence, AlphaTrans translates and executes the test fragments on the recomposed translated project. Super column Test Translation shows the results of test translation and execution. Column TNEF indicates the ratio of AMFs where execution of translated tests never reached them. Columns ATP through ATF show the number of AMFs that AlphaTrans executed using translated test fragments, categorized per the test execution results. For $2.88\%$ of AMFs, all the test fragments that covered them were marked as pass (ATP). For $3.72\%$ and $5.44\%$, at least one (OTF) or more than one test (MTF) failed. All the test fragments failed for $2.62\%$ of them. Note that these numbers (TNEF, ATP, OTF, MTF, and ATF) add up to $56.57\%$ of AMFs covered by source program tests.
@@
For cases with test failure, columns RE and AF show the breakdown of whether test failure was due to assertion failure or runtime error. We can observe that most test executions terminated with a runtime error due to translation bugs and never reached the assert statement. Our manual investigation confirms that a high rate of runtime errors is due to a relatively small number of fragments with translation bugs. Although test decomposition helps with test translation coupling effect2), there is still a high degree of runtime errors due to the long call chains in these projects (the average number of methods executed per test in the original and decomposed test suites are $27.4$ and $21.8$). As a result, the overall pass rate, i.e., the percentage of recomposed test fragments for the translated projects that pass (column TPR), is low. These results also confirm the necessity of in-isolation testing through language interoperability by AlphaTrans.
@@
We also calculated the number of AMFs that GraalVM could not execute (numbers under GE column) but translated test fragments exercised (column M1). All indicates the number of AMFs with all passing tests, including test fragment with assert statement, indicating the validation of functional equivalence. Some corresponds to the number of AMFs with at least one passing test, which indicates runtime validation. Overall, test translation can validate the functional correctness and runtime behavior of $30$ and $118$ fragments that GraalVM could not exercise.
@@
Finally, we analyzed translations using PyLint (pyl, 2024), which scores Python files on a scale of 00 to $10$ based on how Pythonic the code is. All AlphaTrans translations achieved scores of $10$, mainly because (1) LLMs inherently generate idiomatic code and (2) AlphaTrans uses Black (bla, 2024) for formatting the translations extracted from the LLM response. These results confirm that the translations are all Pythonic, i.e., they adhere to coding standards and best practices.
@@
Summary. AlphaTrans effectively performs compositional translation and validation of $17,874$ fragments, achieving overall $96.40\%$ syntactical correctness ($98.80\%$ for AMFs), $27.03\%$ runtime behavior validation (GS+M1 Some), and $25.14\%$ functional equivalence (GS+M1 All).

### v5

@@ 5.1. Type Translation
Automatically resolving types is a challenging problem (Terekhov and Verhoef, 2000; Guizzo et al., 2023), and a large body of work has attempted to address this, mostly using symbolic rule-based approaches (Qiu, 1999; Goerge et al., 2010; Broggi and Liu, 2023; Roziere et al., 2020; Eman J. Coco and Osman, 2018; Lano and Siala, 2024). AlphaTrans employs a Retrieval-Augment Generation (RAG) (Lewis et al., 2020) technique for finding equivalent types in the target language. To that end, it first extracts all the types in the source language of a given project. Custom application types are resolved during the translation as AlphaTrans translates fragments and classes in the target language. For the remaining types, it crawls the API documentation of the source language and retrieves the relevant description of each type.
To form the prompt, [^3] Due to space limit, we omit the prompt; please refer to our artifacts (Lab, 2025a) to see the prompts used for type resolution. AlphaTrans uses the retrieved description and instructs the model with an in-context example to return the most relevant type in the target language, given the use of types in the source language and the retrieved description. To account for potential hallucination in LLM’s response, i.e., returning a type that does not exist in Python, AlphaTrans employs a simple Python script, uses the translated type as an annotation, executes the script, and keeps the ones that have no syntactic or runtime issue. The types in the source language and their corresponding in the target language form a data structure called universal type mapping. In practice, AlphaTrans reuses or augments the mapping when translating new projects.
@@ 5.2. Skeleton Construction
At the first step, AlphaTrans creates Python classes corresponding to each application class in Java. The fields in the classes are set to None, and AlphaTrans uses the information in schema (§4.2.1) to ensure the naming corresponds to the type of their access modifier in the source language. In the example of Figure 4, the translation of field private Option option; in Java is __option: Option = None. The classes also include method signatures, with their body set to pass. AlphaTrans uses the universal type mapping (§5.1) to create relevant types in the method signature. Once the initial skeleton is created, AlphaTrans leverages the extracted call graph during program decomposition to detect circular dependencies. If such dependencies exist, AlphaTrans resolves them with local imports. For inner classes, AlphaTrans unfolds them in Python and uses dot notation to access specific methods and fields (e.g., Class.methodName). Finally, AlphaTrans implements best practices in Python and subclasses all abstract classes and interfaces from abc.ABC class. Here, ABC is a class from the abc module in the Python standard library, which is used for defining abstract base classes.
@@ 6. Compositional Translation and Validation
AlphaTrans translates method fragments in reverse call order. It takes the project’s call graph, removes the back edges to make it acyclic, computes topological order (i.e., linear ordering of its vertices), and translates method fragments (corresponding to vertices) in reverse topological order. Field fragments are independent; therefore, AlphaTrans translates them before method fragments. The algorithm takes a fragment $F$, LLM $M$, Skeleton $S,$ and a series of parameters as inputs, translates the fragment, recomposes the skeleton with the successfully translated fragment, and returns translation outcomes: (1) syntax check (denoted by the “non-parseable” and “parseable” labels), (2) functional equivalence check (denoted by the “graal-fail”, “graal-success”, and “graal-error” labels), and (3) translated test execution check (denoted by “not-exercised”, “test-fail”, and “test-success” labels).
AlphaTrans employs iterative and feedback-based prompting. That is, if one of the mentioned checks fails, e.g., the translated fragment is not syntactically correct, it prompts the model for another translation. To control the number of iterations, AlphaTrans uses a reprompting budget (repromptBudget). The algorithm takes the minimum ($min_{budget}$) and maximum ($max_{budget}$) values for the budget and dynamically sets reprompting budget to a number between the lower and upper bounds based on coverage information, e.g., the budget is close to $max_{budget}$ for a fragment if it is exercised multiple times (high hit rate based on coverage information) by different unit tests. The rationale is to give more importance to fragments covered by more tests to eventually increase translation validation success.
@@ 44       $break$;
[…]
Finally, it translates and executes decomposed fragment tests: if there are no eligible tests (a test becomes eligible if all its dependencies are translated) for the fragment, AlphaTrans simply assigns the “not-exercised” label to the fragment and moves on to the next one (lines 21–23). Otherwise, it translates the tests, executes them to validate the fragment, and assigns test outcome labels in TVO (lines 24–31). In case of a test failure, AlphaTrans extracts all involved fragments and reprompts them with feedback extracted from test execution.
Due to inherent intra- and inter-procedural dependencies in real-world projects, the number of fragments involved in reprompting could be high, logarithmically increasing the translation time. AlphaTrans filters out those with GraalVM label “graal-success”, ranks the remaining based on suspiciousness score, and reprompts $topK$ suspicious fragments. The suspiciousness score for fragments is calculated such that a fragment with more failing tests will get a higher score and, therefore, ranked higher among other fragments.
Our prompt template consists of five distinct parts (Figure 5). The first part is the persona message used by DeepSeek-Coder-33b-Instruct during instruction fine-tuning and is required for producing the best outputs. The next part introduces the In-Context Learning (ICL) example, which reflects the complexities of code translation and instructs LLM on how to deal with them. The green part shows the natural language instruction given to the model. After describing the objective, the prompt embeds the source Java code along with partial translation as a skeleton, which includes all dependencies and translations of the fragments invoked by the current one. The prompt concludes with ### Response: keyword to guide the model for code generation.
@@ 6.1. Language Interoperability
GraalVM (Würthinger et al., 2013; Oracle, 2024) is a Java Development Kit by Oracle. It offers the Polyglot API (GraalVM, 2024), which allows the integration of programs written in different guest languages within a Java-based host application. In the context of this work, GraalVM allows execution of Python code from Java and vice versa (Abid et al., 2024). AlphaTrans leverages the Polyglot API to perform in-isolation validation of the fragments by replacing the Java implementation of a method with its translated Python version while keeping the rest of the project in Java. It then executes the Java tests covering the fragment to validate the functional equivalence of the translation.
The Polyglot API allows access to Python data objects from Java and vice-versa, as these objects reside in a shared memory space. However, objects must be cast to appropriate types for passing parameters to, and processing returned values from, polyglot calls. The Polyglot API can perform this casting implicitly for only a few simple data types. AlphaTrans builds on top of the Polyglot API to provide a framework to create a Python program state that is isomorphic to the Java program state. The Python translation is restricted to this isomorphic state, and the states are synchronized after method calls to preserve the isomorphism. AlphaTrans allows for the casting of user-defined types as well as several built-in and library types. Using both static and dynamic type information of Java objects, AlphaTrans can disambiguate target types when casting Python objects to Java types. It further preserves object identities and aliasing during such casting and propagates exceptions across language boundaries.
@@ 6.3. Test Translation
Similar to translating application fragments, AlphaTrans also translates test fragments. Using the dependency information captured during static analysis, it crafts prompts for unit tests along with their dependencies for the model to translate. The ICL examples used for prompting test fragments differ from prompts used for translating application method fragments. The focus of ICL examples here is to prevent the LLM from hallucinating the used assert statements in the source PL to the target PL. To construct ICL examples for test fragment translation, we created a pool of in-context examples, where each example shows the Python assert statements equivalent to Java assert statements in the context of a test. When prompting a test fragment, AlphaTrans detects the assert statement in the fragment and retrieves the corresponding examples from the pool. For translated tests, only syntactic validation is performed as there is no other means of validating their translations.
@@ 1. RQ1:
    Effectiveness of AlphaTrans. To what extent AlphaTrans can automatically resolve types from source to target PL? Can AlphaTrans effectively translate real-world projects?
@@ 7.1. Experiment Setup
AlphaTrans uses CodeQL (GitHub, 2024) and tree-sitter (Tree-Sitter, 2024) for static analysis. For running tests, validating translation, and computing coverage, AlphaTrans uses GraalVM $21.0.3+7.1$ (Oracle, 2024), JUnit 4 and 5 (Team, 2024b), Pytest 8.2.1 (pytest dev, 2024), JaCoCo (Team, 2024a), and Python’s coverage (Batchelder, 2024). AlphaTrans works with API- and open-access LLMs. We considered the following criteria for selecting the LLM: (1) for better reproducibility, we prioritized open-access models; (2) due to computing constraints, we wanted an LLM with moderate size ($>10B$ but $<70B$ parameters); (3) the model should perform reasonably well in code-related tasks; and (4) the model should have fast inference time due to the huge number of prompts. Per the mentioned criteria, we selected DeepSeek-Coder-33b-Instruct (Guo et al., 2024) for the main experiments (RQ1–RQ4). We also used GPT-4o, one of the best-performing commercial models, for RQ5 (§7.6) to demonstrate the impact of stronger models on improving the performance of AlphaTrans. We prompted models with the temperature set to 00 for reproducibility and used their default settings for other parameters. For the base prompting (Algorithm 2), we set the minimum and maximum values of the reprompting budget to $3$ and $5$. For the feedback prompting, we set the reprompting budget to $1$, i.e., AlphaTrans attempts to fix issues with feedback only once.
@@ 7.2.1. Effectiveness in Type Resolution and Skeleton Validation:
AlphaTrans extracted $1,797$ distinct types from the source projects and attempted to translate them to equivalent Python types. Of these, $915$ are application types (i.e., classes defined within the Java projects) and were directly resolved during skeleton construction. AlphaTrans prompts DeepSeek-Coder-33b-Instruct to resolve the remaining $882$ and successfully translated $738~{}((915+738)/1,797=91.99\%)$ of them: the generated types passed the syntactic and runtime check. The column ATR in Table 1 shows the results of automated type resolution. Because type resolution is essential to project skeleton construction, we manually checked the type mappings generated by DeepSeek-Coder-33b-Instruct and also translated the $144$ unresolved types.
Through manual investigation of the automatically resolved types, we observed that DeepSeek-Coder-33b-Instruct’s type resolution for $182$ cases, while correct, can be improved. For example, AlphaTrans translated java.io.File, a class concerning file manipulation functionality to str. The resolved type can represent file paths in Python but lacks features for file manipulation. We suspect this translation is impacted by the Java use case provided in the prompt. While this translation is correct for the use case, we replaced it with pathlib.Path to have a more generic type mapping. We also augmented the type mapping with additional types in the target language for $38$ types. For example, AlphaTrans translated java.nio.Buffer to bytearray, which is correct as they both provide a mutable sequence of bytes with efficient in-place modifications. However, array.array and memoryview also provide similar functionality with efficient and low-level data manipulation capabilities. Consequently, we augmented type mapping to typing.Union[bytearray, array.array, memoryview]. Given that type mapping can be reused, this one-time manual effort increases the chance of AlphaTrans’s success on unseen projects.
@@
Using the universal type mapping, AlphaTrans successfully creates and validates project skeletons in the target PL, achieving $100\%$ syntax and runtime validation (column SV in Table 1). The skeleton validation step ensures all module imports, class structures, method signatures, and type annotations are done properly, making the subsequent steps easier. Applying AlphaTrans to unseen projects, if a class skeleton cannot be validated, AlphaTrans removes it from the target project, updates the skeleton based on the class dependencies, and proceeds to the next phase.
@@ 7.2.2. Effectiveness in Compositional Translation and Validation:
Table 2 shows the compositional translation and validation results. The AMF column indicates the total number of application method fragments. The numbers in subsequent columns demonstrate the effectiveness of AlphaTrans in the translation and validation of AMFs only. [^4] Please refer to our artifact (Lab, 2025a) for details of all the translation and validation of other fragments shown in Table 1. The Syntax Check column indicates the percentage of AMFs that pass syntactic validation. Column SNEF shows the percentage of AMFs not covered by source project tests. AlphaTrans successfully generates syntactically correct code ($98.80\%$ of AMFs and $96.40\%$ of all fragments—field, application, and test fragments—across ten subjects). We also observe that $43.43\%$ of AMFs are not covered during the execution of any test, i.e., we cannot go beyond syntactic check and validate their runtime behavior or functional equivalence.
For $56.57\%$ of the AMFs that are covered by source project tests, AlphaTrans attempts to validate their functional equivalence using GraalVM. Multi-column GraalVM shows the percentage of AMFs that GraalVM executes and successfully validates (GS), executes but there is a test assertion failure (GF), and cannot execute due to its limitation (GE) mentioned in §6.1. $24.50\%$ (min=$6.01\%$ and max=$70.70\%$), $16.23\%$ (min=$1.04\%$ and max=$32.50\%$), and $15.84\%$ (min=$3.13\%$ and max=$25.11\%$) of AMFs resulted in Graal Success, Graal Fail, and Graal Error, respectively. Note that these numbers add up to $56.57\%$ of AMFs that were covered by source project tests. With respect to only covered methods, GraalVM Success is $43.30\%$. Furthermore, our analysis shows that a high portion of methods that are not covered by tests are either abstract methods or getter/setter methods. If their translations are syntactically correct, they are also likely functionally equivalent, which can ramp up the success rate (§7.5). Spearman Rank Order Correlation (Spearman, 1904) indicates a strong positive correlation between method coverage (Table 1) and GS numbers ($\rho=0.92$), confirming that with better method coverage, validated AMFs are very likely to be higher.
@@
Regardless of GraalVM’s validation for functional equivalence, AlphaTrans translates and executes the test fragments on the recomposed translated project. The Test Translation multi-column shows the results of test translation and execution. Column TNEF indicates the ratio of AMFs where execution of translated tests never reached them. Columns ATP through ATF show the number of AMFs that AlphaTrans executed using translated test fragments, categorized per the test execution results. For $2.88\%$ of the AMFs, all the test fragments that covered them were marked as pass (ATP). For $3.72\%$ and $5.44\%$, at least one test (OTF) or more than one test (MTF) failed. All the test fragments failed for $2.62\%$ of the AMFs. Note that these numbers (TNEF, ATP, OTF, MTF, and ATF) add up to $56.57\%$ of the AMFs that are covered by source project tests.
@@
For cases with test failure, columns RE and AF show the breakdown of whether test failure was due to assertion failure or runtime error. We can observe that most test executions terminated with a runtime error due to translation bugs and never reached an assert statement. Our manual investigation confirms that a high rate of runtime errors is due to a relatively small number of fragments with translation bugs. Although test decomposition helps with test translation coupling effect2), there is still a high degree of runtime errors due to long call chains in these projects (the average number of methods executed per test in the original and decomposed test suites are $27.4$ and $21.8$, respectively). As a result, the overall pass rate, i.e., the percentage of recomposed test fragments for the translated projects that pass (column TPR), is low. These results also confirm the necessity of in-isolation testing through language interoperability by AlphaTrans.
@@
We also calculated the number of AMFs that GraalVM could not execute (numbers under GE column) but translated test fragments exercised (column M1). All indicates the number of AMFs with all passing tests, including test fragments with assert statements, indicating the validation of functional equivalence (with respect to the source project tests). Some corresponds to the number of AMFs with at least one passing test, which indicates runtime validation. Overall, test translation validates the functional correctness and runtime behavior of $30$ and $118$ fragments that GraalVM could not exercise.
@@
Finally, we analyzed translations using PyLint (pyl, 2024), which scores Python files on a scale of 00 to $10$ based on how Pythonic the code is. All AlphaTrans translations achieved scores of $10$, mainly because (1) LLMs inherently generate idiomatic code and (2) AlphaTrans uses Black (bla, 2024) for formatting the translations extracted from the LLM response. These results confirm that the translations are all Pythonic, i.e., they adhere to Python coding standards and best practices.
@@
Summary. AlphaTrans effectively performs compositional translation and validation of $17,874$ fragments, achieving overall $96.40\%$ syntactic correctness ($98.80\%$ for AMFs), $27.03\%$ runtime behavior validation (GS+M1 Some), and $25.14\%$ functional equivalence (GS+M1 All).